# Hesitation and Tolerance in Recommender Systems

### v1

@@ 5.2.1. Datasets
Taobao is an e-commerce platform offering millions of items for sale across various categories. User interactions with items on the platform typically involve behaviors such as clicking, adding to cart, or adding to favorites before making a purchase. As discussed in Section 1, clicking on an item signals exploration, but it does not necessarily reflect the user’s definitive interest in the item. On an e-commerce platform like Taobao, a purchase is a stronger indication of a user’s interest in the item. Similarly, adding an item to the cart or favorites also strongly suggests interest, as it indicates an intent to purchase. In this study, we define clicking without any subsequent actions (e.g., adding to cart, favorites, or purchasing), as negative feedback from the user. We term this behavior as “tolerance behavior”. We clarify that tolerance is fundamentally different from not clicking on a product. When a user chooses not to click, they demonstrate disinterest without investing time. In contrast, tolerance behavior occurs when a user spends time exploring a product in search of potential interest but ultimately feels that their time was wasted, potentially diminishing their overall experience.
@@
Kuaishou is a video-sharing mobile app that allows users to watch short videos on various topics. However, like Taobao, simply clicking on or watching a video doesn’t strongly indicate a user’s interests. A better measure is the time spent watching, a longer watch time usually shows greater interest. This watch time also depends on users’ browsing habits (e.g., patient vs impatient users) and video length (e.g., a few seconds vs a few minutes). We define “tolerance behavior” as the ratio of a user’s watch time to the video’s duration. If a user watches a video with a browsing ratio below their average, it is considered tolerance behavior.
@@
Dataset Reference Week Investigation Week #User #Item
Taobao $1$ $7$, Jun $8$ $14$, Jun $143,475$ $185,315$
$15$ $21$, Sep $22$ $28$, Sep $143,514$ $203,057$
Kuaishou $9$ $15$, Apr $16$ $22$, Apr $26,480$ $5,429,071$
$23$ $29$, Apr $30$ Apr - $6$ May $26,885$ $6,001,034$
5.2.2. Experimental Setting
@@
5.2.3. Tolerance Behavior vs Engagement
In Figure 5, we categorize users based on their tolerance-related behaviors and present the proportion of users within each group who exhibit a decrease in engagement. Specifically, on Taobao, users are grouped by the number of items they clicked on without taking any subsequent action, while on Kuaishou, they are grouped by their average watch ratio during the reference week.
On Taobao, a higher number of items clicked without subsequent actions within a week correlates with increasing tolerance behaviors and a more significant reduction in user engagement. This pattern indicates that when users spend more time on items that do not result in a purchase, their overall activity on the platform tends to decline. Conversely, users who exhibit fewer than 10 tolerance behaviors in a week are more likely to remain active on the platform. Similarly, on Kuaishou, a reduction in the browsing ratio is associated with increase in user tolerance. Users are less inclined to engage with videos they find unappealing. As tolerance levels rise, a growing proportion of users demonstrate reduced activity on Kuaishou, potentially leading to user attrition.
Figure 5 also illustrates that the trend of decreased user engagement with increasing tolerance is consistently observed across different time frames. This finding suggests that tolerance levels significantly influence user satisfaction and retention rates on these platforms. Hence, although tolerance behavior is associated with exploratory actions, e.g., clicks on Taobao or video watching on Kuaishou, it does not necessarily reflect definitive interest. Instead, it provides a weaker signal of user preference on the items. Users often show initial interest and an intention to explore items at first glance but may not take further actions, such as making a purchase or continuing to watch. This indicates that only certain features of the items appeal to them, while others fail to sustain their engagement.
6. Tolerance Behavior for User Retention
The findings from the questionnaire, as presented in Section 4, emphasize a critical insight: tolerance behavior towards recommended items negatively impacts user loyalty to the platform. This underscores the importance of addressing the tolerance signal as a means to optimize recommender systems. By effectively managing this signal, platforms can enhance user retention, which is an essential business metric (Jannach and Jugovac, 2019). In this section, we detail our experimental approach to integrating tolerance behavior into the recommender systems, aiming to improve their performance.
@@ 6.1. Instantiation of Tolerance Behavior
Conventionally, the objective function of a recommendation model is to optimize the following binary cross-entropy loss:
@@
where $\mathcal{Y}$ denotes the positive user-item interaction set and $\mathcal{Y}^{-}$ denotes the negative set. Specifically, $(u,i)$ refers to a user-item pair where user $u$ has interacted with item $i$, such as by clicking on it. Conversely, $(u,j)$ indicates a user-item pair where user $u$ has not interacted with item $j$, and $y$ is the ground-truth label and $\hat{y}$ represents the prediction of recommendation model.
@@
(3) $\mathcal{L}_{T}=-\sum_{(u,i)\in\mathcal{Y}^{P}}y_{ui}\cdot\log\hat{y}_{ui}-\beta\times\sum_{(u,k)\in\mathcal{Y}^{T}}y_{uk}\cdot\log\hat{y}_{uk}-\sum_{(u,j)\in\mathcal{Y}^{-}}(1-y_{uj})\cdot\log(1-\hat{y}_{uj}).$
where $\beta<1.0$ serves as the discounting and down-scaling hyperparameter of tolerance signals. In our evaluation, $\beta$ is dynamically determined based on the user’s content viewing completion rate against the average user completion rate, and then normalized to the interval $[0,1]$. Notably, the primary difference between Equations 1 and 3 is that tolerance signals are separated from the original positive samples, allowing them to be treated as “weak positive samples”.
@@ 6.2. Online Experiments on User Retention
Figure 6. Experiment platform, Youku App, an online video platform with tens of millions of active users.
@@ 6.2.1. Experiment Setting
We partner with the Youku App video platform, which has tens of millions of active users, to evaluate our algorithm. The experiments are conducted on the double-column recommendation list on the second screen of the app’s homepage, as illustrated in Figure 6. We slightly refine the platform’s ranking model (SIM) (Pi et al., 2020) by adjusting the criteria for positive samples based on our findings. Previously, a simple click counted as a positive sample. Now, based on our user study, a positive sample requires a click accompanied by a viewing duration exceeding the user’s average viewing percentage. If a user clicks but watches for less time than their average, we label this as a tolerance sample. Situations with no click remain classified as negative samples.
@@ 6.2.2. Main Results
Table 5 summarizes the A/B test results, where $\mathcal{R}$ represents the change in retention rate, i.e., the increase in the proportion of users who remain active after being presented with recommended content. The results indicate that adjusting the labels consistently improves the Day-2 retention rate across both independent experimental sessions. This outcome underscores the potential of integrating nuanced behavioral signals into recommender systems to capture fine-grained user preferences and engagement, thereby improving the user retention. It also highlights that more sophisticated behavioral modeling increases user satisfaction and loyalty.
Recent studies have explored the user retention problem in other platforms (Zhu et al., 2024; Ding et al., 2023b; Cai et al., 2023). For instance, Cai et al. (2023) modeled user retention optimization as a Markov decision process and developed a retention critic learning, which achieves the retention improvements by around $0.053\%\sim 0.063\%$ in the online A/B test over video recommendation scenario. Ding et al. (2023b) optimized user retention from a contrastive multi-instance learning perspective, which obtains $0.65\%$ next-day user retention improvement over content recommendation scenario. Despite different online testing platforms, our strategy significantly improves average 7-day user retention with subtle model modifications or interventions, reaching $0.67\%$ and $0.36\%$ increase in two separate online tests. These results represent a significant improvement for practical business applications given their large user base, demonstrating our solution’s robustness and effectiveness.
@@
2024-04-20 $+0.794\%$ 2024-04-28 $+0.670\%$ 2024-08-21 $+0.27\%$ 2024-09-15 $+1.62\%$
2024-04-21 $-0.120\%$ 2024-04-29 $-0.694\%$ 2024-08-22 $+0.14\%$ 2024-09-16 $+0.13\%$
2024-04-22 $+1.368\%$ 2024-04-30 $+1.695\%$ 2024-08-23 $-1.82\%$ 2024-09-17 $+1.42\%$
2024-04-23 $-0.604\%$ 2024-05-01 $+0.576\%$ 2024-08-24 $+0.45\%$ 2024-09-18 $+0.33\%$
2024-04-24 $+0.094\%$ 2024-05-02 $-0.132\%$ 2024-08-25 $-0.59\%$ 2024-09-19 $-0.88\%$
2024-04-25 $+0.991\%$ 2024-05-03 $+0.045\%$ 2024-08-26 $+0.14\%$ 2024-09-20 $+2.12\%$
2024-04-26 $+1.915\%$ 2024-05-04 $+0.417\%$ 2024-08-27 $+2.70\%$ 2024-09-21 $+0.72\%$
Average $+0.670\%$ Average $+0.360\%$ Average $+0.15\%$ Average $+0.58\%$
@@
2024-04-20 $-1.92\%$ 2024-04-28 $-0.79\%$ 2024-08-21 $+2.08\%$ 2024-09-15 $-2.16\%$
2024-04-21 $-1.54\%$ 2024-04-29 $+1.19\%$ 2024-08-22 $-1.02\%$ 2024-09-16 $-2.95\%$
2024-04-22 $-3.04\%$ 2024-04-30 $-4.63\%$ 2024-08-23 $+0.00\%$ 2024-09-17 $+0.17\%$
2024-04-23 $+1.02\%$ 2024-05-01 $+1.33\%$ 2024-08-24 $-0.98\%$ 2024-09-18 $-3.32\%$
2024-04-24 $+1.37\%$ 2024-05-02 $-1.48\%$ 2024-08-25 $+0.00\%$ 2024-09-19 $-3.99\%$
2024-04-25 $-0.88\%$ 2024-05-03 $-1.81\%$ 2024-08-26 $+3.30\%$ 2024-09-20 $-0.48\%$
2024-04-26 $-3.10\%$ 2024-05-04 $-2.77\%$ 2024-08-27 $-3.19\%$ 2024-09-21 $+0.35\%$
Average $-1.18\%$ Average $-1.28\%$ Average $+0.03\%$ Average $-1.77\%$
@@ 6.2.3. Impact on Dwell Time
The observed increase in user retention suggests improved loyalty, yet the concurrent decrease in dwell time (see Table 6) reveals critical insights into engagement strategies. While traditional recommender systems often prioritize prolonged user sessions, our findings suggest that simply extending session length does not guarantee a superior experience. Instead, the shorter dwell times may reflect the removal of hesitation or indecision, which contributes little to overall satisfaction or perceived value.
This finding highlights that the quality of engagement is more important than the duration. Long sessions do not necessarily foster loyalty unless users are satisfied with the interactions. Recommender systems should focus on delivering value efficiently, prioritizing content that resonates with users over artificially inflating session length with moments of uncertainty or dissatisfaction. In practice, this means addressing subtle behavioral cues like hesitation or tolerance behaviors. By enhancing genuine user experiences and meeting their needs more effectively, platforms can build deeper loyalty and long-term engagement, paving the way for sustainable, user-focused recommendation strategies.

### v2

@@ 4.1. Datasets
Taobao is an e-commerce platform offering millions of items for sale across various categories. User interactions with items on the platform typically involve behaviors such as clicking, adding to cart, or adding to favorites before making a purchase. As discussed in Section 1, clicking on an item signals exploration, but it does not necessarily reflect the user’s definitive interest in the item. On an e-commerce platform like Taobao, a purchase is a stronger indication of a user’s interest in the item. Similarly, adding an item to the cart or favorites also strongly suggests interest, as it indicates an intent to purchase. In this study, we define clicking without any subsequent actions (e.g., adding to cart, favorites, or purchasing), as negative feedback from the user. We term this behavior as “tolerance behavior”. We clarify that tolerance is fundamentally different from not clicking. When a user chooses not to click, they demonstrate disinterest without investing time. In contrast, tolerance behavior occurs when a user spends time exploring a product in search of potential interest but ultimately feels that their time was wasted, potentially diminishing their overall experience.
@@
Kuaishou is a video-sharing mobile app that allows users to watch short videos on various topics. However, like Taobao, simply clicking on or watching a video doesn’t strongly indicate a user’s interests. A better measure is the time spent watching; a longer watch time usually shows greater interest. This watch time also depends on users’ browsing habits (e.g., patient vs impatient users) and video length (e.g., a few seconds vs a few minutes). We define “tolerance behavior” as the ratio of a user’s watch time to the video’s duration. If a user watches a video with a browsing ratio below their average, it is considered tolerance behavior.
@@
Dataset Ref. Week Investigation Week #User #Item
Taobao 11 - $7$, Jun 88 - $14$, Jun $143,475$ $185,315$
$15$ - $21$, Sep $22$ - $28$, Sep $143,514$ $203,057$
Kuaishou 99 - $15$, Apr $16$ - $22$, Apr $26,480$ $5,429,071$
$23$ - $29$, Apr $30$ Apr - $6$ May $26,885$ $6,001,034$
4.2. Experimental Setting
@@
These proxies are user-normalized and reflect systematic under-engagement relative to a user’s typical behavior; they therefore distinguish tolerance from random noise or low-quality items. Their consistency across users and their strong predictive value for reduced future engagement further support their interpretation as tolerance signals.
@@
4.3. Tolerance Behavior vs. Engagement
Figure 4 shows how tolerance behaviors map onto changes in user engagement. On Taobao, people who frequently clicked on products but took no further action were less likely to stay active over time. By contrast, those with relatively few such tolerance events continued to engage more consistently. On Kuaishou, a similar trend emerged: when users watched videos for less than their typical viewing ratio, their overall activity declined. Across both platforms, the pattern is clear that higher tolerance is strongly linked to reduced engagement. These findings suggest that while tolerance may begin with exploratory actions (e.g., clicking or briefly watching), it rarely translates into sustained interest. Instead, tolerance marks a fragile state of attention: users are curious enough to look, but not convinced enough to continue. Over time, this gap erodes satisfaction and increases the likelihood of leaving the platform.
Finding 4.
Tolerance is a weak signal of interest, capturing curiosity without commitment, and its accumulation predicts disengagement and eventual attrition.
5. Online Experiment: Tolerance Behavior for User Retention
Based on the findings from our surveys (Section 3.2) highlight a critical insight: tolerance behavior toward recommended items undermines user loyalty. This makes it essential to treat tolerance signals not as noise but as actionable feedback for optimizing recommender systems.
However, analyses of offline datasets confirm that tolerance behaviors correlate with reduced engagement, but engagement alone only reflects declining interest. Without directly measuring retention, it remains unclear whether modeling tolerance can truly enhance user loyalty. To overcome this limitation, we conduct large-scale online A/B tests on a commercial video platform. By explicitly incorporating tolerance signals into the ranking model and evaluating their impact on next-day retention, we provide causal evidence of their practical value.
Through online experiments, we address our third research question: can modeling tolerance improve retention in practice? By demonstrating causal improvements in retention, a key business metric (Jannach and Jugovac, 2019), our study establishes tolerance modeling as an effective and scalable strategy for enhancing recommender performance.
@@ 5. Online Experiment: Tolerance Behavior for User Retention
The online A/B experiments were reviewed and approved through both our internal process and the platform’s ethics committee. All data used were strictly de-identified and aggregated, ensuring that no personally identifiable information was accessed. In line with platform policies, users could opt out of the influence of experimental recommendation algorithms at any time.
@@ 5.1. Instantiation of Tolerance Behavior
Conventionally, the objective function of a recommendation model is to optimize the binary cross-entropy loss as:
@@
where $\mathcal{Y}$ denotes the positive user-item interaction set and $\mathcal{Y}^{-}$ denotes the negative set. Specifically, $(u,i)$ refers to a user-item pair where user $u$ has interacted with item $i$, such as click on it; $(u,j)$ indicates a user-item pair where user $u$ has not interacted with item $j$, and $y$ is the ground-truth and $\hat{y}$ represents the prediction of recommendation model.
@@
(3) $\mathcal{L}_{T}=$ $-\sum_{(u,i)\in\mathcal{Y}^{P}}y_{ui}\cdot\log\hat{y}_{ui}-\beta\times\sum_{(u,k)\in\mathcal{Y}^{T}}y_{uk}\cdot\log\hat{y}_{uk}$
$-\sum_{(u,j)\in\mathcal{Y}^{-}}(1-y_{uj})\cdot\log(1-\hat{y}_{uj}),$
where $\beta\in(0,1)$. Here, $\beta$ is a down-weighting coefficient derived directly from how much a user’s viewing duration falls below their own historical completion baseline, after normalization to the interval $(0,1)$. Thus, $\beta$ reflects the degree of under-engagement relative to the user’s typical behavior, with larger deviations resulting in smaller $\beta$ values. Note that $\beta$ is not a tunable hyperparameter but a deterministic quantity computed from observed behavior; cases with zero viewing duration are treated directly as negative samples and therefore do not produce $\beta$ values. The key difference between Equations 1 and 3 is that tolerance signals are separated from conventional positives and assigned discounted weights $\beta$, allowing them to be modeled as “weak positive” signals. These formulations acknowledge that not all clicks are equal: some represent genuine interest, while others capture hesitation or regret.
@@ 5.2. Online Experiments on User Retention
Figure 5. Experiment platform, a Chinese online video and streaming service platform with tens of millions of active users.
@@ 5.2.1. Experiment Setting
We partner with a Chinese online video and streaming service platform, which has tens of millions of active users, to evaluate our algorithm. The experiments are conducted on the double-column recommendation list on the second screen of the app’s homepage, as illustrated in Figure 5. We slightly refine the platform’s ranking model (SIM) (Pi et al., 2020) by adjusting the criteria for positive samples based on our findings. Previously, a simple click counted as a positive sample. Now, based on our user study, a positive sample requires a click accompanied by a viewing duration exceeding the user’s average viewing percentage. If a user clicks but watches for less time than their average, we label this as a tolerance sample. Situations with no click remain classified as negative samples.
@@ 5.2.2. Main Results
Table 5 summarizes the A/B test results, where $\mathcal{R}$ represents the change in retention rate, i.e., the increase in the proportion of users who remain active after being presented with recommended content. The results indicate that adjusting the labels consistently improves the Day-2 retention rate across both independent experimental sessions. This outcome underscores the potential of integrating nuanced behavioral signals into recommender systems to capture fine-grained user preferences and engagement, thereby improving the user retention. It also highlights that more sophisticated behavioral modeling increases user satisfaction and loyalty. Recent studies have explored the user retention problem in other platforms (Zhu et al., 2024; Ding et al., 2023; Cai et al., 2023). For instance, Cai et al. (2023) modeled user retention optimization as a Markov decision process and developed a retention critic learning, which achieves the retention improvements by around $0.053\%\sim 0.063\%$ in the online A/B test over video recommendation scenario. Ding et al. (2023) optimized user retention from a contrastive multi-instance learning perspective, which obtains $0.65\%$ next-day user retention improvement over content recommendation scenario. Despite different online testing platforms, our strategy significantly improves average 7-day user retention with subtle model modifications or interventions, reaching $0.67\%$ and $0.36\%$ increase in two separate online tests. These results represent a significant improvement for practical business applications given their large user base, demonstrating our solution’s robustness and effectiveness.
@@
2024-04-20 $0.794\%$ 2024-04-28 $0.670\%$ 2024-08-21 $0.27\%$ 2024-09-15 $1.62\%$
2024-04-21 $0.120\%$ 2024-04-29 $0.694\%$ 2024-08-22 $0.14\%$ 2024-09-16 $0.13\%$
2024-04-22 $1.368\%$ 2024-04-30 $1.695\%$ 2024-08-23 $1.82\%$ 2024-09-17 $1.42\%$
2024-04-23 $0.604\%$ 2024-05-01 $0.576\%$ 2024-08-24 $0.45\%$ 2024-09-18 $0.33\%$
2024-04-24 $0.094\%$ 2024-05-02 $0.132\%$ 2024-08-25 $0.59\%$ 2024-09-19 $0.88\%$
2024-04-25 $0.991\%$ 2024-05-03 $0.045\%$ 2024-08-26 $0.14\%$ 2024-09-20 $2.12\%$
2024-04-26 $1.915\%$ 2024-05-04 $0.417\%$ 2024-08-27 $2.70\%$ 2024-09-21 $0.72\%$
Average $0.670\%$ Average $0.360\%$ Average $0.15\%$ Average $0.58\%$
@@
2024-04-20 $1.92\%$ 2024-04-28 $0.79\%$ 2024-08-21 $2.08\%$ 2024-09-15 $2.16\%$
2024-04-21 $1.54\%$ 2024-04-29 $1.19\%$ 2024-08-22 $1.02\%$ 2024-09-16 $2.95\%$
2024-04-22 $3.04\%$ 2024-04-30 $4.63\%$ 2024-08-23 $0.00\%$ 2024-09-17 $0.17\%$
2024-04-23 $1.02\%$ 2024-05-01 $1.33\%$ 2024-08-24 $0.98\%$ 2024-09-18 $3.32\%$
2024-04-24 $1.37\%$ 2024-05-02 $1.48\%$ 2024-08-25 $0.00\%$ 2024-09-19 $3.99\%$
2024-04-25 $0.88\%$ 2024-05-03 $1.81\%$ 2024-08-26 $3.30\%$ 2024-09-20 $0.48\%$
2024-04-26 $3.10\%$ 2024-05-04 $2.77\%$ 2024-08-27 $3.19\%$ 2024-09-21 $0.35\%$
Average $1.18\%$ Average $1.28\%$ Average $0.03\%$ Average $1.77\%$
@@ 5.2.3. Impact on Dwell Time
The observed gain in retention points to stronger loyalty, yet the accompanying drop in dwell time (see Table 6) offers a different perspective on engagement. Traditional recommender systems often equate success with longer sessions, but our findings suggest that duration alone does not guarantee satisfaction. In fact, shorter sessions may reflect the removal of hesitation or indecision — moments that add length without adding value. This shift underscores that the quality of engagement matters more than its duration. Long sessions only foster loyalty when users feel their time is well spent. Recommender systems should therefore aim to deliver value efficiently, prioritizing content that resonates rather than prolonging use through uncertainty or frustration. In practice, this means attending to subtle cues such as hesitation and tolerance, and adapting quickly when they appear. By focusing on interactions that minimize wasted effort and maximize relevance, platforms can build deeper trust and long-term engagement, advancing toward more sustainable, user-centered recommendation strategies.
6. Discussion
Our study identifies hesitation and tolerance as two intermediate states in user–recommender interactions that are often overlooked when engagement is reduced to clicks and views. Across surveys, behavioral traces, and an online experiment, we find that hesitation is nearly ubiquitous and frequently leads to tolerance, which correlates with frustration, diminished trust, and reduced satisfaction. We now discuss these findings in relation to broader theories of user interaction and intelligent systems, outlining how our conceptualization extends current understandings of how users navigate and interpret recommendations. We then describe how these interaction states challenge long-standing recommender-system assumptions and metrics, followed by limitations and future work.
6.1. User Interaction States in HCI
Although we started from a recommender system perspective, this interaction state modeling connects to long-standing theories of human action, decision-making, and interpretation. For example, Norman’s framework of gulfs of execution and evaluation (Whitenton, 2018) helps explain hesitation and tolerance. Hesitation reflects a widened gulf of execution, where users deliberate because it is unclear how or whether to act on a recommendation. Tolerance reflects a gulf of evaluation, where users persist with unwanted content but struggle to assess or act upon their dissatisfaction. Both states reveal frictions that are not captured by binary accept/reject models. While related concepts appear in prior work, this study is the first to explicitly model hesitation and tolerance in the context of recommender systems—an increasingly important environment where such frictions often go unexamined.
Our findings also speak to questions of agency and autonomy raised in Value Sensitive Design (Friedman, 1996; Friedman et al., 2013). Tolerance exemplifies a loss of agency: users continue with unwanted recommendations because exit paths are unclear or costly. Hesitation, meanwhile, illustrates increased cognitive load, as users must invest additional effort to resolve uncertainty. Recognizing these states connects recommender system evaluation to broader theories of effort, engagement, and satisfaction.
Much of our earlier focus emphasized business-oriented metrics such as clicks, dwell time, or retention. Here, we take a step back to foreground the human side of these interactions. By situating hesitation and tolerance within HCI theory, we highlight how user experience, agency, and wellbeing can be compromised when recommendations are designed only for profit. This shift places human value at the center, showing that evaluation must move beyond transactional outcomes to include experiential quality, fairness, and trust — dimensions that are central to HCI and critical in an era where recommendations mediate so much of everyday life.
Much of our earlier focus emphasized business-oriented metrics such as clicks, dwell time, or retention. Here, we take a step back to foreground the human side of these interactions. By situating hesitation and tolerance within established theories of user experience, agency, and wellbeing, we highlight how these states reveal compromises that arise when recommendations are designed primarily for engagement maximization.
This shift places human value at the center, showing that evaluation must move beyond transactional outcomes to include experiential quality, fairness, and trust — dimensions that are critical in an era where recommendations shape so much of everyday life.
6.2. Rethink Clicks and CTR: What Exactly Has Been “Disrupted”?
Click-through rate (CTR) has long been treated as the default yardstick for recommender systems. Its popularity stems from convenience, clicks are plentiful, easy to log, and appear to represent interest. For years, both research benchmarks and industry KPIs have revolved around this single number. Yet our findings complicate this picture. By surfacing the states of hesitation and tolerance, we show that clicks often capture uncertainty or even dissatisfaction rather than genuine interest. What looks like “positive engagement” in the logs may actually be the residue of doubt, wasted time, or frustration, experiences that erode trust and weaken retention.
This calls into question the assumption that more clicks or longer watch times are always better. Instead, recommender systems should be optimized for what we term low-regret satisfaction: interactions that respect users’ time, reduce unnecessary effort, and deliver value that feels meaningful. In other words, engagement metrics must evolve from counting actions to capturing the quality of experience those actions represent.
6.3. Implication: For Researchers, Designers, Developers, and Users
6.3.1. For Researchers: Metrics and Methodologies
Our findings suggest that evaluation needs to account for the process of decision-making, not just its outcomes. A click should not be treated as a uniform positive signal; rather, it sits within a broader spectrum that includes hesitation, shallow exploration, and tolerance-driven disengagement. By introducing conceptual “layers” between pre-click and post-click, we can distinguish strong positives from weak or costly ones. In our video experiments, for example, shallow viewing, defined relative to a person’s own viewing baseline, proved to be a reliable indicator of tolerance and produced measurable improvements when modeled explicitly.
Moving forward, we argue that evaluation should adopt metrics that make hidden costs and user burdens visible, rather than relying solely on raw clicks or dwell time. To this end, we propose three new measures:
  • Time-to-Disinterest (TTD). The elapsed time from exposure to an item until the user explicitly rejects it or exits. TTD captures how long it takes for users to recognize irrelevance, highlighting inefficiencies in the recommendation process.
  • Hesitation Tax. The additional time cost a user incurs within a session before concluding that they are not interested. This metric surfaces the “wasted effort” dimension of hesitation, turning a subjective frustration into a measurable burden.
  • Regret-Adjusted CTR (rCTR). A click-through rate adjusted by tolerance weights, discounting clicks that lead to shallow engagement or rapid abandonment. rCTR aims to separate curiosity-driven clicks from genuine interest, producing a more faithful indicator of user value.
Together, these metrics move evaluation beyond surface-level activity counts, toward measures that acknowledge regret, wasted effort, and erosion of trust.
Equally important is a shift toward long-term objectives. Retention and sustained activity provide a more faithful picture of system impact than immediate clicks or watch time. Our experiments show that even small adjustments, such as reclassifying shallow views, can lead to meaningful gains in next-day retention. This underscores that success should not be defined by more clicks or longer sessions, but by cultivating efficient, low-regret interactions that accumulate into durable value over time.
6.3.2. For Designers: Interaction and Feedback Mechanisms
Our findings point to several design directions that can help reduce tolerance and protect user trust.
Make rejection effortless when designing the interactions. Participants consistently emphasized the frustration of wasted time. Providing quick-access options such as “Not relevant,” “Clickbait,” or “Duplicate” within one tap lowers the cost of signaling dissatisfaction. This prevents tolerance sessions from being misread as positive signals and gives users a sense of control.
Try to recognize lightweight disengagement. Everyday actions like fast-forwarding, speeding up playback, skipping ahead, or exiting early should be interpreted as meaningful weak negatives rather than ignored or rolled up into “watch time.” Treating these subtle signals as part of the feedback loop aligns system interpretation with actual user intent.
Let the recommendation session fail fast, adapt quickly. When tolerance signals appear, the system should pivot, offering alternative content or adjusting its strategy, rather than nudging users to finish what they already find unappealing. Our online experiments confirmed that shortening “invalid watch time” does not reduce retention; in fact, it helps maintain satisfaction by respecting users’ time. Designing for tolerance means valuing efficiency and agency: letting users say “no” easily, treating small disengagements as signals, and responding swiftly when recommendations miss the mark.
6.3.3. For Developers: Modeling and Engineering
Our results highlight several opportunities for developers to refine how recommender systems interpret user feedback.
For modeling, move beyond binary labels. Instead of treating every click as a positive signal, developers can adopt a richer classification scheme that separates strong positives (deep viewing or follow-up actions), tolerance (shallow or unfulfilled engagement), and clear negatives (no interaction). Modeling these distinctions allows systems to capture subtle but important differences in user intent.
For user understanding, respect individual baselines. Whether an interaction counts as tolerance depends heavily on personal habits and content length. A completion rate that looks “shallow” for one user may be normal for another. Using relative thresholds — benchmarked against each individual’s own history — offers a more faithful representation of behavior.
Digging more data for modeling, like enhancing logging and observability. To operationalize hesitation and tolerance, platforms need to log more than clicks and dwell time. Events such as fast-forwarding, lingering on a detail page without action, or adjusting playback speed provide valuable signals that can feed into feature stores and model training. Capturing these nuances at scale opens the door to recommendation models that better align with real user experiences.
Engineering for tolerance means treating feedback as multi-layered, personalized, and observable — bridging raw interaction logs with the complexity of human decision-making.
6.3.4. For Users: Agency and Well-being
Finally, our findings highlight the importance of giving users a stronger voice in shaping their recommendations. Many participants expressed frustration when systems misinterpreted exploratory actions as genuine interest, leading to repeated exposure to unwanted content.
Offer users the right to explanation. A simple, one-click option such as “This is not what I expected” can empower users to clarify their intent. By treating this form of regret-driven feedback as especially meaningful, systems can quickly correct course and avoid reinforcing negative loops. Supporting such mechanisms helps break the chain from tolerance to fatigue to eventual churn. More importantly, it signals respect for users’ time and agency, making the recommender system feel less like a black box and more like a partner that listens and adapts.
User-facing controls that acknowledge regret and allow quick clarification build trust, reduce frustration, and prevent tolerance from escalating into abandonment.
6.4. Limitations for this work
Our work is not without limitations. First, the scenarios we examined are narrow: our surveys focused on e-commerce and short-video platforms, and our online validation was restricted to one specific placement in a single video app. How tolerance plays out in other domains, such as news, live streaming, or productivity tools, remains uncertain. Second, both our samples and contexts carry potential biases. Survey participants were primarily recruited from developer communities, which may not reflect the diversity of everyday users. Similarly, the online experiments reflect the design choices of one platform and may not generalize elsewhere. Third, our evaluation lens was limited. We focused mainly on Day-2 retention and aggregate watch time, leaving out subjective measures such as satisfaction, perceived regret, or trust, as well as longer-term engagement indicators.
Finally, modeling tolerance touches on ethical questions. Specifically, inferring “negative emotions” from behavioral traces risks over-interpretation. Transparency and user agency are essential if such signals are to be used responsibly, ensuring that such modeling empowers rather than exploits users.
6.5. Future Directions
There are several promising avenues for extending this work. Larger-scale surveys across more diverse populations and platforms could validate whether our findings generalize beyond the current sample. Beyond breadth, deeper studies are needed to unpack the psychological underpinnings of hesitation, uncertainty, curiosity, and exploratory intent, and to design interventions that are both precise and respectful.
Longer-term experiments, such as rolling A/B tests over weeks or months, would shed light on how tolerance accumulates and affects sustained retention. Cross-domain studies, spanning news feeds, long-form video, live streaming, and multi-stage e-commerce, would test the boundaries of where tolerance matters most.
On the modeling side, future work could integrate tolerance into reinforcement learning or multi-objective optimization, not just as another feature but as a way of shaping rewards around low-regret engagement. Comparative studies against purchase-driven ranking or dwell-time-based methods would help clarify its added value.
Another promising direction is to investigate UI-level signals and cues that help users disambiguate interest more efficiently, thereby reducing the likelihood that hesitation escalates into tolerance. Integrating such user interface and experience considerations with model-level tolerance handling could offer a more holistic approach to reducing user regret.
Finally, personalization remains a critical challenge. Adaptive thresholds that reflect individual habits, combined with qualitative approaches such as interviews and field studies, would ground the notion of a “hesitation tax” in lived experience. Building shared datasets or benchmarks with tolerance annotations could further accelerate progress, inviting the research community to move beyond clicks and toward metrics that respect time, effort, and trust.